OpenStudy (anonymous):

Hockey team receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 64 points. They won 11 more games than they tied. How many wins and how many ties did the team have? How many wins did the team have?

OpenStudy (anonymous):

number of wins = x number of ties = y

OpenStudy (anonymous):

2x + y = 64 x = y+11 solve them

OpenStudy (anonymous):

sub 2nd into 1st 2(y+11) +y = 64 3y +22 = 64 y= 14 sub into the 2nd eqn

OpenStudy (anonymous):

x= 14 +11 = 25

OpenStudy (anonymous):

therefore 25winss

OpenStudy (anonymous):

thanks.....;)

OpenStudy (anonymous):

how many did they tied?

OpenStudy (anonymous):

well, if you look back at the working out should be easy

OpenStudy (anonymous):

14